### v5

# Electromagnetic Inverse Scattering from a Single Transmitter

Abstract
Electromagnetic Inverse Scattering Problems (EISP) seek to reconstruct relative permittivity from scattered fields and are fundamental to applications like medical imaging. This inverse process is inherently ill-posed and highly nonlinear, making it particularly challenging, especially under sparse transmitter setups, e.g., with only one transmitter. While recent machine learning-based approaches have shown promising results, they often rely on time-consuming, case-specific optimization and perform poorly under sparse transmitter setups. To address these limitations, we revisit EISP from a data-driven perspective. The scarcity of transmitters leads to an insufficient amount of measured data, which fails to capture adequate physical information for stable inversion. Accordingly, we propose a fully end-to-end and data-driven framework that predicts the relative permittivity of scatterers from measured fields, leveraging data distribution priors to compensate for the incomplete information from sparse measurements. This design enables data-driven training and feed-forward prediction of relative permittivity while maintaining strong robustness to transmitter sparsity. Extensive experiments show that our method outperforms state-of-the-art approaches in reconstruction accuracy and robustness. Notably, we demonstrate, for the first time, high-quality reconstruction from a single transmitter. This work advances practical electromagnetic imaging by providing a new, cost-effective paradigm to inverse scattering.
Equal contribution, random order.
Correspondence: {ez_cheng, tianchunxun, ou524u}@stu.pku.edu.cn
{wtzhu, maxiaoxuan, yizhou.wang}@pku.edu.cn
Project page: https://gomenei.github.io/SingleTX-EISP/
1 Introduction
Electromagnetic waves can penetrate object surfaces, making them essential for non-invasive imaging (Geng et al., 2024; O’Loughlin et al., 2018). At the core of electromagnetic imaging lies the Electromagnetic Inverse Scattering Problems (EISP), which seeks to reconstruct an object’s relative permittivity from measured scattered electromagnetic field (Nikolova, 2011). By solving EISP, we can accurately recover internal structures without physical intrusion (Song et al., 2005), enabling a range of scientific and industrial applications, such as safer and more cost-effective alternatives to X-rays and MRI scans (Bevacqua et al., 2021; O’Loughlin et al., 2018; Nikolova, 2011). Typically, EISP necessitate a large number of transmitters and receivers to acquire sufficient measurement data. This requirement, however, leads to increased operational time and higher costs, thereby limiting the practical applicability of electromagnetic imaging techniques (Leigsnering et al., 2011). In contrast, reducing the number of transmitters offers significant advantages, including lower costs and easier deployment in constrained environments.(Baraniuk and Steeghs, 2007; Anitori et al., 2010)
Figure 1: Comparison between our method and the previous state-of-the-art. Left: Img-Interiors (Luo et al., 2024) requires case-specific optimization to reconstruct the permittivity. In contrast, our method is a data-driven framework that operates in an end-to-end, feed-forward manner for solving inverse scattering. Right: Our method yields more accurate reconstructions than Img-Interiors (Luo et al., 2024). It remains robust even with a single transmitter and achieves real-time inference with over $70,000\times$ speed-up.
However, the inherent ill-posed nature of EISP poses significant challenges to accurate reconstruction (Pan et al., 2011; Chen, 2018; Li et al., 2019; Zhong et al., 2016; Luo et al., 2024), particularly when only a limited number of transmitters are available. The scarcity of transmitters leads to an insufficient amount of measured data, which fails to capture adequate physical information for stable inversion. As a result, approaches relying solely on physical mechanisms(Slaney et al., 1984; Belkebir et al., 2005; Chen, 2009; Zhong and Chen, 2011) often fail to achieve accurate reconstruction. Conventional numerical methods such as backpropagation (BP) (Belkebir et al., 2005), generally fail to produce reliable reconstructions under such limited-data conditions. Recent machine learning-based approaches like PGAN (Song et al., 2021) and Physics-Net (Liu et al., 2022) often start with an initial solution derived from numerical methods, i.e., BP, and frame the problem as an image-to-image translation task. With only a limited number of transmitters available, reliance on BP becomes a critical bottleneck. When BP fails, these methods are unable to correct its errors, as they are not fully end-to-end, ultimately leading to inaccurate reconstructions. The most recent method Img-Interiors (Luo et al., 2024) integrates physical mechanisms into neural networks and performs case-by-case optimization. However, in limited-transmitter scenarios, even after optimization has converged, the resulting reconstructions may still diverge substantially from the ground truth (Fig.˜2), underscoring the intrinsic ambiguity of the inverse problem.
To address these limitations, we propose a fully end-to-end and data-driven framework that predicts the relative permittivity of scatterers from measured fields, leveraging data distribution priors to compensate for the scarcity of observational data. Specifically, our model takes the measured fields and the spatial coordinate of a position as input and directly predicts the relative permittivity at that location using Multilayer Perceptron (MLP)s, and is trained in a fully end-to-end manner against the ground-truth data. Our approach bypasses traditional numerical methods like BP, thereby avoiding the inherent constraints associated with conventional inversion techniques in limited-transmitter scenarios and fully exploiting the advantages of data-driven learning. This simple yet effective design enables efficient training across datasets and supports fast, feed-forward inference to achieve accurate and stable reconstruction predictions.
Extensive experiments demonstrate that our method outperforms existing State-of-the-Art (SOTA) methods on multiple benchmark datasets, especially under the challenging single-transmitter setting, where all previous methods fail (Fig.˜5). It generalizes well to diverse scenarios and can be naturally extended to 3D scenes while maintaining high reconstruction accuracy. In summary, our contributions are threefold:
1) We systematically analyze the difficulty of lacking physical information faced by EISP in the setting of few transmitters, and point out that the missing information can be supplemented by data distribution priors.
2) Based on our analysis, we propose a fully end-to-end and data-driven model that does not rely on traditional numerical methods.
3) Extensive experiments show that our method outperforms existing SOTA approaches, especially under the challenging single-transmitter setting, marking a concrete step toward cost-effective and practical electromagnetic imaging solutions.
2 Related Work
2.1 Electromagnetic Inverse Scattering Problems (EISP)
Solving EISP is to determine the relative permittivity of the scatterers based on the scattered field measured by the receivers, thereby obtaining internal imaging of the object. The primary challenges of EISP arise from its nonlinearity, ill-posedness, and errors introduced by the discretization (Pan et al., 2011; Chen, 2018; Li et al., 2019; Zhong et al., 2016; Luo et al., 2024). Traditional methods for solving EISP can be categorized into non-iterative (Slaney et al., 1984; Devaney, 1981; Habashy et al., 1993; Belkebir et al., 2005) and iterative (Chen, 2009; Zhong and Chen, 2011; Xu et al., 2017; Habashy et al., 1994; van den Berg et al., 1999) approaches. Non-iterative methods, such as the Born approximation (Slaney et al., 1984), the Eytov approximation (Devaney, 1981; Habashy et al., 1993), and the BP method (Belkebir et al., 2005), solve nonlinear equations through linear approximations, which inevitably lead to poor quality of the results. For better reconstruction quality, iterative methods (Zhong and Chen, 2009; Chen, 2009; Zhong and Chen, 2011; Xu et al., 2017; Habashy et al., 1994; van den Berg et al., 1999; Gao et al., 2015) such as 2-fold Subspace Optimization Method (SOM) (Zhong and Chen, 2009) and Gs SOM (Chen, 2009) are proposed. To further overcome the ill-posedness of EISP, diverse regularization approaches and prior information have been widely applied (Oliveri et al., 2017; Shen et al., 2014; Liu et al., 2018; Anselmi et al., 2018). However, all of these methods are not generalizable and can be time-consuming because of the iterative schemes (Liu et al., 2022).
2.2 Machine Learning for EISP
Recent studies shift to leverage neural networks to solve this problem and demonstrate promising results (Geng et al., 2021; Li et al., 2024). Some work (Wei and Chen, 2019; Li et al., 2019; Zhang et al., 2020; Xu et al., 2021; Sanghvi et al., 2019; Song et al., 2021; Liu et al., 2022) adopt a two-stage strategy: they use non-iterative methods such as BP (Belkebir et al., 2005) to generate initial estimates, which are then refined using image-to-image neural networks. While these approaches offer a degree of generalization, they are not end-to-end and remain dependent on BP initialization (Belkebir et al., 2005), which becomes their bottleneck. When physical data are too insufficient to reconstruct the scatterer, especially under single-transmitter settings, these approaches tend to “hallucinate” outputs according to unreliable initialization rather than predict the scatterer based on measured field (see Fig.˜5). A more recent approach, Img-Interiors (Luo et al., 2024), integrates scattering mechanisms into the network architecture and achieves accurate reconstructions. However, it requires case-specific optimization, limiting generalization and making it vulnerable to local minima, often leading to failure in complex settings (see Figs.˜4 and 6). Moreover, it fails under a single transmitter setting even when the optimization may have already converged because of ambiguity. While our method is also learning-based, it is an end-to-end feed-forward framework that simultaneously achieves generalization through data-driven learning. As a result, it consistently outperforms SOTA methods, particularly in the challenging single-transmitter setting where previous approaches fail.
3 Revisiting EISP
In this section, we revisit EISP and uncover its fundamental challenge: the inherent ill-posedness stemming from information scarcity.
Preliminary. In the forward process, the transmitters produce incident electromagnetic field $\mathbf{E}^{\text{i}}$ to the scatterer, generating scattered electromagnetic field $\mathbf{E}^{\text{s}}$. EISP is the inverse problem of the forward process. That is, for an unknown scatterer, we apply certain incident field $\mathbf{E}^{\text{i}}$ to it, and measure the scattered field $\mathbf{E}^{\text{s}}$ as our input. Our goal is to reconstruct the relative permittivity $\bm{\epsilon}_{r}$ throughout the scatterer. For a detailed background introduction of the physical model of EISP, please refer to our supplementary material (Appx.˜B).
Specifically, the incident field $\mathbf{E}^{\text{i}}$ excites the induced current $\mathbf{J}$. Using the method of moments (Peterson et al., 1998), the total field $\mathbf{E}^{\text{t}}$ for a given transmitter can be expressed as (Colton and Kress, 2013):
$\mathbf{E}^{\text{t}}=\mathbf{E}^{\text{i}}+\mathbf{G}^{\text{d}}\cdot\mathbf{J},$ (1)
where $\mathbf{E}^{\text{t}}$ is a vector of length $M^{2}$, and $\mathbf{G}^{\text{d}}$ is a constant $M^{2}\times M^{2}$ matrix representing the discrete free-space Green’s function in $\mathcal{D}$. The induced current field $\mathbf{J}$ satisfies:
$\mathbf{J}=\operatorname{Diag}(\bm{\xi})\cdot\mathbf{E}^{\text{t}},$ (2)
where $\bm{\xi}=\bm{\epsilon}_{r}-1$, $\operatorname{Diag}(\bm{\xi})$ represents a diagonal matrix whose leading diagonal consists of $\bm{\xi}$. Then $\mathbf{J}$ serves as a new source to emit $\mathbf{E}^{\text{s}}$. For $N_{r}$ receivers, the scattered field $\mathbf{E}^{\text{s}}$ can be got through $\mathbf{E}^{\text{s}}=\mathbf{G}^{\text{s}}\cdot\mathbf{J}$, where $\mathbf{G}^{\text{s}}$ is a constant $N_{r}\times M^{2}$ matrix representing the discrete Green’s function. Since $N_{r}\ll M^{2}$ in practice, reconstructing the induced current $\mathbf{J}$ from the scattered field $\mathbf{E}^{\text{s}}$ is ill-posed.
Figure 2: Difficulties that previous methods faced under a single-transmitter setting. (a) BP cannot reconstruct the scatterer. (b) Physics-Net makes incorrect guesses. (c) Although the reconstuction result of Img-Interiors is consistent with the measured field, the reconstructed scatterer itself is completely different from the ground truth.
Reduction of measured data. EISP is fundamentally challenged by nonlinearity and ill-posedness, especially when the amount of measured data is significantly reduced, such as under single-transmitter settings. We divide previous work into three categories and systematically analyze the difficulties they faced under this setting. (a) Conventional numerical approaches, such as BP (Belkebir et al., 2005), employ linear approximations, which limit their reconstruction quality. As shown in Fig.˜2, BP cannot even reconstruct a rough shape of the scatterer. (b) Machine learning methods based on conventional numerical approaches Song et al. (2021); Liu et al. (2022), such as Physics-Net. Although Physics-Net can leverage data-driven training to compensate for missing physical information, its strong dependency on BP initialization becomes a critical bottleneck. When BP fails, the model cannot correct the error of BP because it is not fully end-to-end, resulting inaccurate reconstructions, as shown in Fig.˜2. (c) Machine learning methods based on implicit functions, such as Img-Interiors (Luo et al., 2024). Img-Interiors reconstructs a scatterer through case-by-case optimization. As shown in Fig.˜2, we use the scatterer reconstructed by Img-Interiors to simulate the scattered field, and the result closely matches the measured field. However, the scatterer itself deviates significantly from the ground truth, which shows the intrinsic ambiguity of the inverse problem. The core conclusion is that the severe information deficit makes a direct solution to the inverse problem fundamentally intractable. Consequently, any such attempt is bound to be fragile, highlighting the need for an alternative paradigm.
4 Method
4.1 Overview
To address the aforementioned limitations, we introduce our end-to-end, data-driven framework for EISP, as illustrated in Fig.˜3. Our method employs an MLP that takes space coordinates $\bf{x}$} and corresponding scattered field measurements $\mathbf{E}^{\text{s}}$ as input, and directly outputs the relative permittivity $\bm{\epsilon}_{r}$ at the specified locations. This approach effectively learns the mapping between scattered field $\mathbf{E}^{\text{s}}$ and relative permittivity $\bm{\epsilon}_{r}$ through training on diverse scattering scenarios, thereby incorporating essential data distribution priors to compensate for the lack of physical information caused by insufficient measurements.
In the following, we detail our model architecture (Section˜4.2), and the training losses (Section˜4.3).
4.2 Model Architecture
Based on the forward formulation of EISP in Sec.˜3, where the scattered field measurements serve as input and the relative permittivity distribution represents the output, we design an end-to-end learning framework that directly learns this complex nonlinear mapping. As illustrated in Fig.˜3, our approach employs an MLP that approximates the inverse mapping from spatial coordinates and scattered field data to the relative permittivity values, formulated as:
$\hat{\bm{\epsilon}_{r}}(\mathbf{x}_{\text{i}})=F_{\theta}(\mathbf{E}^{\text{s}},\gamma(\mathbf{x}_{\text{i}})),\mathbf{x}_{\text{i}}\in\mathbb{R}^{2},$ (3)
where $\mathbf{x}_{\text{i}}$ represents the spatial coordinate, $\mathbf{E}^{\text{s}}$ denotes the scattered field measured by all receivers, $F_{\theta}(\cdot)$ is an MLP with trainable parameters, and $\hat{\bm{\epsilon}_{r}}(\mathbf{x}_{\text{i}})$ is the predicted relative permittivity at the corresponding position. Recall that in Sec.˜3, for a single transmitter, the scattered field $\mathbf{E}^{\text{s}}$ is discretized as a real-valued vector of dimension $2N_{r}$, containing the real and imaginary parts of the measurements from all $N_{r}$ receivers. In the multiple transmitter configuration, $\mathbf{E}^{\text{s}}$ is constructed by combining the complex measurement data from all $N$ transmitters, resulting in a real-valued vector of dimension $2N\cdot N_{r}$ that represents the wave propagation and scattering behavior under diverse illumination conditions provided by transmitters at different locations.
Figure 3: Overview of our method. Our pipeline is built around a MLP that serves as the inverse solver. Given the scattered field measurements $\mathbf{E}^{\text{s}}$ from all transmitters and receivers, along with a spatial query $\bf{x}$}, the MLP directly predicts thee relative permittivity $\hat{\bm{\epsilon}_{r}}({\bf{x}})$. To enhance spatial expressiveness, we apply positional encoding $\gamma({\bf{x}})$ to the query position. During training, dashed lines indicate the supervision signals applied.
To enhance the model’s capacity to represent high-frequency features, we apply positional encoding to the spatial coordinates $\mathbf{x}_{\text{i}}$, mapping them into a higher-dimensional Fourier feature space using the encoding function: $\gamma(x)=[\sin(x),\cos(x),\ldots,\sin(2^{\Omega-1}x),\cos(2^{\Omega-1}x)]$, where the hyperparameter $\Omega$ controls the spectral bandwidth.
The complete relative permittivity distribution $\hat{\bm{\epsilon}_{r}}$ is reconstructed by sampling the MLP at all grid points $\{\mathbf{x}_{\text{i}}\}_{i=1}^{M^{2}}$: $\hat{\bm{\epsilon}_{r}}=\{F_{\theta}(\mathbf{E}^{\text{s}},\gamma(\mathbf{x}_{\text{i}})\}_{\text{i}=1}^{M^{2}}.$
4.3 Training
Our training objective is defined by a single loss function designed to directly supervise the reconstruction accuracy of the relative permittivity distribution. The loss is formulated as: $\mathcal{L}=\|\hat{\bm{\epsilon}_{r}}-\bm{\epsilon}_{r}\|^{2}.$ where $\hat{\bm{\epsilon}_{r}}$ denotes the predicted relative permittivity and $\bm{\epsilon}_{r}$ represents the ground truth. By minimizing this Mean Squared Error (MSE) loss between the predicted and true permittivity values, the model learns to infer the material properties directly from the scattered field measurements, effectively leveraging the data distribution priors to overcome the ill-posedness of the inverse problem. This simplified loss function ensures stable and efficient training.
5 Experiments
5.1 Setup
Datasets. We train and test our method on standard benchmarks used for EISP following previous work (Wei and Chen, 2019; Song et al., 2021; Liu et al., 2022). To enhance training efficiency and model robustness, datasets that share identical transmitter and receiver configurations are combined into a unified training set. 1) Synthetic Circular-cylinder dataset (Circular(Luo et al., 2024) is synthetically generated comprising images of cylinders with random relative radius, number, and location and permittivity. 2) Synthetic MNIST dataset (MNIST(Deng, 2012) contains grayscale images of handwritten digits. For the two synthetic datasets, we follow previous work(Luo et al., 2024; Deng, 2012), we evaluate two levels of noise: 5% and 30%. 3) Real-world Institut Fresnel’s database (IF(Geffrin et al., 2005) contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel. 4) Synthetic 3D MNIST dataset (3D MNIST(de la Iglesia Castro, ) contains 3D data of handwritten digits. 5) 3D ShapeNet dataset (3D ShapeNet(Wu et al., 2015) contains 3D data of various shapes. For more details about datasets, please refer to Appx.˜C.
Table 1: Quantitative comparison results with SOTA methods. For Circular and MNIST datasets, we report results under two noise levels: 5% and 30%. The best results are shown in bold, and the second-best results are underlined.
 Method MNIST (5%) MNIST (30%) Circular (5%) Circular (30%) IF
MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
  Number of Transmitters: $N=16$ $N=8/18$
BP (2005) 0.177 0.719 16.43 0.178 0.716 16.38 0.052 0.905 27.41 0.053 0.904 27.42 0.190 0.779 16.19
2-fold SOM (2009) 0.154 0.757 20.93 0.156 0.738 20.84 0.031 0.917 32.23 0.038 0.889 30.63 - - -
Gs SOM (2009) 0.072 0.923 28.31 0.081 0.901 27.13 0.023 0.946 35.40 0.024 0.937 34.89 0.184 0.790 17.00
BPS (2019) 0.093 0.909 25.00 0.105 0.891 23.90 0.027 0.963 33.00 0.029 0.956 32.42 0.310 0.664 17.05
PGAN (2021) 0.084 0.916 25.80 0.091 0.910 25.31 0.026 0.966 35.56 0.032 0.947 33.91 0.121 0.926 24.78
Physics-Net (2022) 0.075 0.932 26.17 0.093 0.906 24.58 0.027 0.934 32.72 0.030 0.927 32.08 0.170 0.788 18.48
Img-Interiors (2024) 0.200 0.863 26.41 0.336 0.760 19.01 0.036 0.947 35.05 0.047 0.932 32.62 0.153 0.837 23.26
Ours 0.039 0.978 32.11 0.050 0.966 29.91 0.020 0.965 36.92 0.024 0.954 35.19 0.094 0.916 24.89
  Number of Transmitters: $N=1$
BP (2005) 0.194 0.698 15.40 0.194 0.696 15.40 0.065 0.892 25.30 0.065 0.892 25.30 0.199 0.770 16.29
2-fold SOM (2009) 0.432 0.556 12.49 0.828 0.382 9.45 0.060 0.859 26.63 0.157 0.639 20.07 - - -
Gs SOM (2009) 0.460 0.598 15.31 0.404 0.557 14.91 0.046 0.888 29.62 0.051 0.862 28.77 0.192 0.779 16.66
BPS (2019) 0.189 0.774 18.75 0.205 0.744 17.97 0.045 0.891 29.29 0.055 0.862 27.68 0.348 0.669 16.18
PGAN (2021) 0.133 0.867 21.69 0.153 0.830 20.41 0.033 0.932 32.02 0.040 0.914 29.94 0.248 0.680 16.85
PhysicsNet (2022) 0.137 0.798 19.98 0.152 0.783 19.38 0.055 0.887 26.60 0.056 0.890 26.48 0.175 0.771 17.45
Img-Interiors (2024) 0.305 0.604 16.06 0.467 0.484 12.47 0.096 0.855 26.19 0.153 0.806 20.90 0.305 0.705 17.34
Ours 0.085 0.921 26.09 0.127 0.862 22.56 0.031 0.931 33.18 0.038 0.914 31.38 0.128 0.908 24.19
Baselines and Metrics. We maintain the same settings as in previous studies (Wei and Chen, 2019; Song et al., 2021; Sanghvi et al., 2019) to ensure a fair comparison. We compare our method with three traditional methods and four deep learning-based approaches: 1) BP (Belkebir et al., 2005): A traditional non-iterative inversion algorithm. 2) 2-fold SOM (Zhong and Chen, 2009): A traditional iterative minimization scheme by using SVD decomposition. 3) Gs SOM (Chen, 2009): A traditional subspace-based optimization method by decomposing the operator of Green’s function. 4) BPS (Wei and Chen, 2019): A CNN-based image translation method with an initial guess from the BP algorithm. 5) Physics-Net (Liu et al., 2022): A CNN-based approach that incorporates physical phenomena during training. 6) PGAN (Song et al., 2021): A CNN-based approach using a generative adversarial network. 7) Img-Interiors (Liu et al., 2022): An implicit approach optimized by forward calculation. Following previous work (Liu et al., 2022), we evaluate the quantitative performance of our method using PSNR (Wang and Bovik, 2009), SSIM (Wang et al., 2004), and Relative Root-Mean-Square Error (MSE(Song et al., 2021).
5.2 Comparison with SOTAs
5.2.1 Multiple Transmitter Evaluation
Figure 4: Qualitative comparison under the multiple-transmitter setting. The results are obtained with $N=16$ transmitters and a noise level of 5%. Colors represent the values of the relative permittivity.
We begin by comparing our method against prior approaches under the multiple-transmitter setting, using both synthetic and real datasets for comprehensive evaluation. As shown in the upper part of Tab.˜1, our method achieves comparable or superior performance to the SOTA in most cases, demonstrating how our end-to-end training framework successfully leverage the data prior across diverse data domains.
In addition, we present a qualitative comparison, as shown in Fig.˜4. Traditional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) are only capable of recovering the coarse shape of the scatterer. BPS (Chen, 2018) produces sharp edges, but the reconstructed shapes are often inaccurate. PGAN (Song et al., 2021) achieves accurate shape recovery, yet introduces noticeable background artifacts. Img-Interiors (Luo et al., 2024) can generate high-quality reconstructions, but occasionally fails due to local optima, as it is based on an iterative optimization process (see the last row). In contrast, our method produces accurate and clean reconstructions across all cases, demonstrating both visual fidelity and robustness.
5.2.2 Single Transmitter Evaluation
Figure 5: Qualitative comparison under the single-transmitter setting. Results are obtained with $N=1$ transmitter and noise level of 5%. Colors represent the values of the relative permittivity.
Furthermore, we investigate a highly challenging and practically important setting that has been largely underexplored in previous work: performing EISP with a minimal number of transmitters. Specifically, we consider the most extreme case, using only a single transmitter. As shown in the lower part of Tab.˜1, our method significantly outperforms all previous approaches across all datasets and noise levels. This remarkable performance under such constrained conditions underscores the efficacy of our end-to-end training framework, which successfully encodes and leverages rich data priors to achieve state-of-the-art results across diverse domains.
To better understand this phenomenon, we present qualitative comparisons in Fig.˜5. Traditional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) produce only blurry reconstructions. Deep learning-based methods like BPS (Chen, 2018), Physics-Net (Liu et al., 2022), and PGAN (Song et al., 2021) tend to “hallucinate” the digit, resulting in wrong shape on the MNIST dataset. Img-Interiors (Luo et al., 2024) fails to capture the fundamental morphology of the scatterer, resulting in structurally inaccurate representations that deviate significantly from the ground truth. Among all the methods, only ours can still produce reasonably accurate reconstructions of the relative permittivity under such an extreme condition.
Figure 6: Qualitative comparison under high noise setting. The results are obtained with $N=16$ transmitters and a noise level of 30%. Colors represent the values of the relative permittivity.
5.3 Ablation Study
Noise Robustness. To simulate real-world sensor noise and related perturbations, we evaluate the robustness of the models by adding noise to the scattered field. Moving beyond simple binary testing, we systematically assess the model’s performance across multiple noise levels ranging from 5% to 30% to examine its behavior in various noisy environments. The quantitative results presented in Tab.˜3 demonstrate that our model exhibits smooth and gradual performance degradation as the noise level increases, maintaining excellent reconstruction capability even under strong noise interference as high as 30%. Qualitative visualizations in Fig.˜6 show that most baseline methods exhibit noticeable artifacts or even complete failure under severe noise conditions, while our method remains robust and preserves the essential structure of the target.
Table 2: Ablation study of noise levels effects on MNIST under the multiple-transmitter setting.
[…]
5% 0.039 0.978 32.11
10% 0.039 0.978 32.18
15% 0.043 0.973 31.30
20% 0.043 0.974 31.34
25% 0.046 0.970 30.59
30% 0.050 0.966 29.91
Table 3: Ablation study on training data size under the mutiple-transmitter setting. Noise levels ($5\%$ and $30\%$) in parentheses.
 Data Size MNIST (5%) MNIST (30%)
[…]
100% 0.039 0.978 32.11 0.050 0.966 29.91
75% 0.043 0.974 31.63 0.059 0.956 28.77
50% 0.048 0.968 30.68 0.068 0.944 27.69
25% 0.064 0.948 28.89 0.101 0.902 25.44
Ablation on Training Data Size. To investigate the dependency of model performance on training data volume, we trained our model on varying scales of data from 100% down to 25% and evaluated them on a complete test set. The quantitative results are presented in Tab.˜3. First, our model demonstrates remarkable data efficiency, maintaining strong performance even when trained on partial datasets. Second, the performance degradation becomes substantially more pronounced under high-noise conditions. The performance penalty for data reduction is markedly severer in high-noise scenarios. This pronounced contrast underscores that sufficient training data is crucial for the model to learn robust features capable of countering strong noise interference.
5.4 Reconstruction on 3D data
Figure 7: Qualitative comparison under the single-transmitter setting for 3D reconstruction on 3D MNIST dataset. The results are obtained with a single transmitter ($N=1$). The voxel colors represent the values of the relative permittivity.
Figure 8: Qualitative comparison under the single-transmitter setting for 3D reconstruction on 3D ShapeNet dataset. The results are obtained with a single transmitter ($N=1$). The voxel colors represent the values of the relative permittivity.
Setup and Metrics. Our method can be naturally extended to 3D scenarios. Following previous work (Luo et al., 2024), we employ the Synthetic 3D MNIST (de la Iglesia Castro, ) and extend to 3D ShapeNet (Wu et al., 2015) for training and testing. For evaluation, we adopt 3D versions of the MSE (Song et al., 2021) and Intersection over Union (IoU) as our metrics. Further details on the datasets are provided in Appx.˜C.
Results. We evaluate our method and Img-Interiors under limited-transmitter settings. Quantitative results demonstrate the superiority of our approach: on 3D MNIST, our method achieves an MSE of 0.120 and IoU of 0.769 with $N=1$ transmitter, significantly outperforming Img-Interiors which obtains an MSE of 0.372 and IoU of 0.094 under the same conditions. With $N=6$ transmitters, our results further improve to MSE of 0.094 and IoU of 0.834. For the more complex 3D ShapeNet dataset under $N=1$ configuration, our method obtains an MSE of 0.064 and IoU of 0.762, showcasing its generalization capability to diverse 3D structures. Fig.˜1 provides a comprehensive comparison of reconstruction quality and runtime between the two methods for both $N=1$ and $N=6$ configurations. Fig.˜7 and Fig.˜8 provide visual comparisons of additional 3D reconstruction results on the 3DMNIST and 3DShapeNet datasets. These results show that our method maintains robustness and generalizes effectively across geometrically complex 3D structures, representing significant progress towards practical applications.
6 Conclusion
In this work, we propose a fully end-to-end data-driven framework for electromagnetic inverse scattering that directly predicts relative permittivity from scattered field measurements. By leveraging data distribution priors to compensate for the lack of physical information, our method demonstrates state-of-the-art reconstruction accuracy and robustness, particularly in challenging single-transmitter scenarios where existing methods fail. This work highlights the potential of data-driven approaches to overcome the ill-posedness of inverse problems and provides a practical path toward cost-effective electromagnetic imaging.
Limitations. While our method effectively handles sparse transmitter settings, it cannot deal with different locations of receivers or transmitters, which remains an important direction for future work.
Appendix A The Use of Large Language Models (LLMs)
We use GPT-4 solely for the purpose of polishing our language of manuscript. This includes improving grammatical accuracy and sentence fluency. LLMs play no significant role in research ideation, methodology design, and experiment execution of this paper.
Appendix B Physical Model of EISP
The research subject of EISP is scatterers. We can describe scatterers with their relative permittivity $\epsilon_{r}({\bf{x}})$. The relative permittivity is a physical quantity determined by the material, and it represents the ability to interact with the electromagnetic field of the material. The relative permittivity of vacuum is $1$ (and the relative permittivity of air is almost $1$), while the relative permittivity of scatterers is over $1$. The relative permittivity of metal material is positive infinity, so metal can shield against electromagnetic waves. Thus, scatterers are not composed of metal.
In EISP, the scatterer is placed within the region of interest $\mathcal{D}$. The transmitters are placed around $\mathcal{D}$, emitting incident electromagnetic field ${E}^{\text{i}}$. The incident field then interacts with the scatterer, exciting the induced current $J$. The induced current can act as secondary radiation sources, emitting scattered field $E^{\text{s}}$. In fact, for a point in the scatterer, it cannot distinguish the incident field and the scattered field at this point. So, it interacts with the sum of the incident field and the scattered field, aka the total field $E^{\text{t}}$. The total field can be described by Lippmann-Schwinger equation (Colton and Kress, 2013) as follows:
$E^{\text{t}}({\bf{x}})={E}^{\text{i}}({\bf{x}})+k_{0}^{2}\int_{\mathcal{D}}g({\bf{x}},{\bf{x^{\prime}}})J({\bf{x^{\prime}}})d{\bf{x^{\prime}}},{\bf{x}}\in\mathcal{D},$ (4)
where $\bf{x}$} and $\bf{x^{\prime}}$} are the spatial coordinates. $k_{0}$ is the wavelength of the electromagnetic wave determined by the frequency. $g({\bf{x}},{\bf{x^{\prime}}})$ is the free space Green’s function, which represents the impact of the induced current $J$ at the point $\bf{x^{\prime}}$} to the total field at the point $\bf{x}$}. ${\bf{x}}\in\mathcal{D}$ indicates that the total field is with the region of interest $\mathcal{D}$. The relationship between the induced current $J$ and the total field $E^{\text{t}}$ can be expressed as follows:
$J({\bf{x}})=\xi({\bf{x}})E^{\text{t}}({\bf{x}}),$ (5)
where $\xi({\bf{x}})=\epsilon_{r}({\bf{x}})-1$.
The induced current generates scattered field, and we can measure the scattered field with receivers around the region of interest $\mathcal{D}$. The scattered field can be expressed as follows:
$E^{\text{s}}=k_{0}^{2}\int_{\mathcal{D}}g({\bf{x}},{\bf{x^{\prime}}})J({\bf{x^{\prime}}})d{\bf{x^{\prime}}},{\bf{x}}\in S,$ (6)
where ${\bf{x}}\in S$ indicates that the scattered field is measured by the receivers at surface $S$ around $\mathcal{D}$.
Since digital analysis only applies to discrete variables (Mitra, 2001; Taflove et al., 2005), we discretize equations Eq.˜4, Eq.˜5, and Eq.˜6. The region of interest $\mathcal{D}$ is discretized into $M\times M$ square subunits, and we use the method of moment (Peterson et al., 1998) to obtain the discrete scattered field $\mathbf{E}^{\text{s}}$ [^1] We use the bold letters to represent discrete variables.. The discrete version of Eq.˜4 is as follows:
$\mathbf{E}^{\text{t}}=\mathbf{E}^{\text{i}}+\mathbf{G}^{\text{d}}\cdot\mathbf{J},$ (7)
where $\mathbf{G}^{\text{d}}$ is the discrete free space Green’s function from points in the region of interest to points in the region of interest, which is a matrix of the shape $M^{2}\times M^{2}$. The discrete version of Eq.˜5 is as follows:
$\mathbf{J}=\operatorname{Diag}(\bm{\xi})\cdot\mathbf{E}^{\text{t}}.$ (8)
And the discrete version of Eq.˜6 is as follows:
$\mathbf{E}^{\text{s}}=\mathbf{G}^{\text{s}}\cdot\mathbf{J},$ (9)
where $\mathbf{G}^{\text{s}}$ is the discrete free space Green’s function from points in the region of interest to the locations of receivers, which is a matrix of the shape $N_{r}\times M^{2}$.
Appendix C Details of Experimental Setup
C.1 Datasets
The datasets utilized in this work are described in detail below. To enhance model robustness and training efficiency, datasets that share identical measurement settings are pooled to form consolidated training sets. We strictly follow the established rules in EISP (Wei and Chen, 2019; Song et al., 2021; Liu et al., 2022) to set the dataset. We train and test our model on standard benchmarks used for EISP.
1) Circular (Luo et al., 2024) is synthetically generated comprising images of cylinders with random relative radius, number, location, and permittivity between $1$ and $1.5$. $10k$ images are generated for training purposes, and $1.2k$ images for testing.
2) MNIST (Deng, 2012) contains grayscale images of handwritten digits. Similar to previous settings (Wei and Chen, 2019; Zhou et al., 2022), we use them to synthesize scatterers with relative permittivity values between $2$ and $2.5$ according to their corresponding pixel values. The entire MNIST training set containing $60k$ images is used for training purposes, while $1.2k$ images from the MNIST test set are randomly selected for testing.
Figure 9: Positions of transmitters and receivers for 2D data. (a) For Circular dataset and MNIST dataset, we set up $N=16$ transmitters and $N_{r}=32$ receivers. All transmitters and receivers are equally placed. (b) For IF dataset, we set up $N=8$ or $18$ transmitters and $N_{r}=241$ receivers. The transmitters are equally placed (not shown in the figure), and the locations of receivers are determined by the transmitter. For all datasets, the green square represents the region of interest $\mathcal{D}$.
Following previous work (Chen, 2018; Zhou et al., 2022) to generate the above two synthetic datasets, we set operating frequency $f=400$ MHz. The region of interest is a square with the size of $2m\times 2m$. We use $16$ transmitters and $32$ receivers equally placed on a circle with a radius of $3m$ ($N_{r}=32$ and $N=16$). The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in LABEL:fig:schematic_synthetic. The data are generated numerically using the method of moments (Peterson et al., 1998) with a $224\times 224$ grid mesh to avoid inverse crime (Colton et al., 1998). To simulate the noise in actual measurement, we add a $5\%$ level of noise to the scattered field $\mathbf{E}^{\text{s}}$ for regular setting.
3) Real-world IF dataset (Geffrin et al., 2005) contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel. $N=8$ for FoamDielExt and FoamDielInt, $N=18$ for FoamTwinDiel, and $N_{r}=241$ for all the cases. The region of interest is a square with the size of $0.15m\times 0.15m$. Transmitters and receivers are placed on a circle with a radius of $1.67m$. The transmitters are placed equally, and the locations of receivers vary for each transmitter. There is no receiver at any position closer than $60^{\circ}$ from the transmitter, and 241 receivers are placed from $60^{\circ}$ to $300^{\circ}$ with a step of $1^{\circ}$ from the location of the transmitter. The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in LABEL:fig:schematic_IF. In the real measurement, there is only one transmitter at a fixed location, and the scatterer rotates to simulate the transmitter to be in different directions. There is a movable receiver sequentially measures the scattered field at 241 different locations. After the measurement, the scatterer rotates by a certain angle for next measurement. The angle is $45^{\circ}$ for FoamDielExt and FoamDielInt because $N_{r}=8$, and $20^{\circ}$ for FoamTwinDiel because $N_{r}=18$. As for operating frequency, all cases are measured under many different frequencies, and we take the frequency $f=5$ GHz. We evaluate these three scenarios for testing, and use the same settings to synthetically generate $10k$ images of cylinders with random number and location for training purposes.
Figure 10: Positions of transmitters and receivers for 3D MNIST dataset. There are $N=40$ transmitters and $N_{r}=160$ receivers. The green cube represents the region of interest $\mathcal{D}$.
4) Synthetic 3D MNIST (de la Iglesia Castro, ) contains 3D images of handwritten digits. 5) Synthetic 3D ShapeNet (Wu et al., 2015) contains 3D images of various shapes. We use these two datasets to synthesize scatterers with relative permittivity value of $2$. $N_{r}=160$ and $N=40$, and the operating frequency $f=400$ MHz. The region of interest is a cube with the size of $2m\times 2m\times 2m$. The transmitters and receivers are placed at a sphere with the radius of $3m$. For the positions of transmitters, the azimuthal angle ranges from $0^{\circ}$ to $315^{\circ}$ with the step of $45^{\circ}$, and the polar angle ranges from $30^{\circ}$ to $150^{\circ}$ with the step of $30^{\circ}$. For the positions of receivers, the azimuthal angle ranges from $0^{\circ}$ to $348.75^{\circ}$ with the step of $11.25^{\circ}$, and the polar angle ranges from $30^{\circ}$ to $150^{\circ}$ with the step of $30^{\circ}$. The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in Fig.˜10. The entire 3D MNIST dataset is used, including $5k$ images for training purposes and $1k$ images for testing. For 3D ShapeNet, we take $5k$ images from $5$ different categories for training purposes and $500$ images for testing.
C.2 Metrics
Following previous work (Liu et al., 2022), we evaluate the quantitative performance of our method using PSNR (Wang and Bovik, 2009), SSIM (Wang et al., 2004), Relative Root-Mean-Square Error (MSE) (Song et al., 2021) and IoU. For PSNR, SSIM and IoU, a higher value indicates better performance. For MSE, a lower value indicates better performance. MSE is a metric widely used in EISP, which is defined as follows:
$\mathrm{MSE}=({\frac{1}{M\times M}\sum_{m=1}^{M}\sum_{n=1}^{M}|\frac{\hat{\bm{\varepsilon}}_{r}(m,n)-\bm{\varepsilon}_{r}(m,n)}{\bm{\varepsilon}_{r}(m,n)}|^{2}})^{\frac{1}{2}},$ (10)
where $\bm{\varepsilon}_{r}(m,n)$ and $\hat{\bm{\varepsilon}}_{r}(m,n)$ are the Ground Truth (GT) and predicted discrete relative permittivity of the unknown scatterers at location $(m,n)$, respectively, and $M\times M$ is the total number of subunits over the Region of Interest (ROI) $\mathcal{D}$.
C.3 Implementation Details
We implement our method using PyTorch. Our MLP consists of 8 layers, each with 512 channels. ReLU activation is used between layers to ensure nonlinear expressiveness. Apply positional encoding to $\bf{x}$} before inputting it into MLP.
During training, we discretize the region of interest $\mathcal{D}$ into a $64\times 64$ grid and optimize the model using the Adam optimizer with default values $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=10^{-8}$, and an initial learning rate of $1\times 10^{-3}$, which remains constant throughout the entire training process. Positional encoding uses frequency $\Omega=10$, and training runs for 200 iterations on an NVIDIA V100 GPU.
Appendix D Additional Experimental Results
This section provides additional experimental results, including statistical validation (Section˜D.1) and extended qualitative comparisons (Section˜D.2). The latter includes more visual comparisons with baseline methods, as well as supplementary ablation studies on noise robustness and training data size.
D.1 Experimental Statistical Significance
Given the inherent stochasticity of the additive noise in our experiments, all quantitative results reported in the main manuscript represent averages over 3 independent trials. The noise samples were independently drawn from a $\mathcal{N}(0,1)$ distribution and subsequently scaled according to both the signal amplitude and the predefined noise ratios (5% and 30%). We report the mean and standard deviation (std) in Table 4, which shows that the experimental variations introduced less than 1% error, confirming the statistical significance of our results.
Table 4: Statistical results of quantitative metrics. For Circular and MNIST datasets, we report the mean and standard deviation (std) over 3 independent trials under two noise levels: 5% and 30%.
  MNIST (5%) MNIST (30%) Circular (5%) Circular (30%)
MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
  Number of Transmitters: $N=16$
mean 0.039 0.978 32.11 0.050 0.966 29.91 0.020 0.965 36.92 0.024 0.954 35.19
std $2.8\times 10^{-5}$ $5.2\times 10^{-5}$ $6.2\times 10^{-3}$ $2.1\times 10^{-4}$ $2.4\times 10^{-4}$ $1.7\times 10^{-2}$ $1.6\times 10^{-5}$ $4.8\times 10^{-5}$ $7.6\times 10^{-3}$ $7.1\times 10^{-5}$ $3.3\times 10^{-4}$ $1.4\times 10^{-2}$
  Number of Transmitters: $N=1$
mean 0.085 0.921 26.09 0.127 0.862 22.56 0.031 0.931 33.18 0.038 0.914 31.38
std $2.1\times 10^{-4}$ $3.1\times 10^{-4}$ $2.0\times 10^{-2}$ $8.1\times 10^{-5}$ $7.2\times 10^{-4}$ $2.2\times 10^{-2}$ $5.8\times 10^{-5}$ $1.3\times 10^{-4}$ $1.2\times 10^{-2}$ $2.6\times 10^{-4}$ $5.2\times 10^{-4}$ $3.3\times 10^{-2}$
D.2 Additional Qualitative Results
Qualitative Comparison.
We present more qualitative comparison on Circular dataset (Luo et al., 2024) and MNIST dataset (Deng, 2012) under settings with different transmitter numbers $N$ and noise levels, as shown in Fig.˜11 to Fig.˜16. We also present qualitative comparison on IF (Geffrin et al., 2005) dataset under the single-transmitter setting, as shown in Fig.˜17. Our method achieves comparable or superior performance to SOTA methods in most cases under multiple-transmitter settings, as Fig.˜11, Fig.˜12, Fig.˜15, and Fig.˜16 indicate. As shown in Fig.˜12 and Fig.˜16, PGAN (Song et al., 2021) introduces noticeable background artifacts due to the lack of consideration of physics. And Fig.˜11 shows that Img-Interiors occasionally fails to converge due to local optima.
For single-transmitter setting ($N=1$), our method significantly outperforms all previous approaches across all datasets, as shown in Fig.˜13 and Fig.˜14. Conventional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) produce only blurry reconstructions. Deep learning-based methods such as BPS (Chen, 2018), Physics-Net (Liu et al., 2022), and PGAN (Song et al., 2021) tend to “hallucinate” the digit, leading to wrong reconstruction on the MNIST dataset. Img-Interiors (Luo et al., 2024) produces results with structural errors that deviate significantly from the true morphology. In contrast, our method can still produce reasonably accurate reconstructions of the relative permittivity under such a challenging condition. This enables practical applications with fewer transmitters, significantly reducing deployment costs while preserving reconstruction quality.
3D Reconstruction.
We present more qualitative comparison on 3D MNIST (de la Iglesia Castro, ) dataset under the single-transmitter setting, as shown in Fig.˜18. Our method successfully approximates permittivity reconstruction even in this difficult setting, whereas Img-Interiors (Luo et al., 2024) fails. Similar to the 2D scenario, Img-Interiors suffers from degeneration and can only produce fixed patterns, failing to capture the shape of scatterers.
Additionally, we evaluate our method on the more complex 3D ShapeNet (Wu et al., 2015) dataset to demonstrate its generalization capability. As shown in Fig.˜19, our approach can successfully reconstruct various complex structures including airplanes, cars and tables under the same single-transmitter configuration, maintaining accurate permittivity distribution recovery. This demonstrates the strong potential of our method for practical real-world applications.
Noise Robustness.
We present qualitative results across multiple noise levels (ranging from 5% to 30%) on the MNIST (Deng, 2012) dataset, providing visual support for the noise robustness analysis. As shown in Fig.˜20, our method maintains high visual fidelity at low noise levels. More importantly, the method maintains stable reconstructions and preserves the essential structure of the scatterer across the entire tested noise range (5% to 30%), demonstrating a smooth and gradual degradation in quality that is fully consistent with the quantitative results.
Ablation on Training Data Size.
We present qualitative results across different training data scales (ranging from 25% to 100%) on the MNIST (Deng, 2012) dataset, offering visual insights into the impact of data volume on reconstruction performance. As shown in Fig.˜21 and Fig.˜22 , our model maintains satisfactory reconstruction integrity even with limited training data.
Figure 11: Qualitative comparison under the multiple-transmitter setting on MNIST dataset. The results are obtained with $N=16$ transmitters and a noise level of 5%. Colors represent the values of the relative permittivity.
Figure 12: Qualitative comparison under the multiple-transmitter setting on Circular dataset. The results are obtained with $N=16$ transmitters and a noise level of 5%. Colors represent the values of the relative permittivity.
Figure 13: Qualitative comparison under the single-transmitter setting on MNIST dataset. The results are obtained with $N=1$ transmitter and a noise level of 5%. Colors represent the values of the relative permittivity.
Figure 14: Qualitative comparison under the single-transmitter setting on Circular dataset. The results are obtained with $N=1$ transmitter and a noise level of 5%. Colors represent the values of the relative permittivity.
Figure 15: Qualitative comparison under high noise setting on MNIST dataset. The results are obtained with $N=16$ transmitters and a noise level of 30%. Colors represent the values of the relative permittivity.
Figure 16: Qualitative comparison under high noise setting on Circular dataset. The results are obtained with $N=16$ transmitters and a noise level of 30%. Colors represent the values of the relative permittivity.
Figure 17: Qualitative comparison under single-transmitter setting on IF dataset. The results are obtained with $N=1$ transmitter, without noise level. Colors represent the values of the relative permittivity.
Figure 18: Qualitative comparison under the single-transmitter setting for 3D reconstruction on 3D MNIST dataset. The results are obtained with single transmitter ($N=1$). The voxel colors represent the values of the relative permittivity.
Figure 19: Qualitative comparison under the single-transmitter setting for 3D reconstruction on 3D ShapeNet dataset. The results are obtained with single transmitter ($N=1$). The voxel colors represent the values of the relative permittivity.
Figure 20: Qualitative results of our noise robustness on the MNIST dataset. The results are obtained with $N=16$ transmitter. Colors represent the values of the relative permittivity.
Figure 21: Qualitative results of training data size ablation on the MNIST dataset. The results are obtained with $N=16$ transmitters and a noise level of 5%.
Figure 22: Qualitative results of training data size ablation on the MNIST dataset. The results are obtained with $N=16$ transmitters and a noise level of 30%.

### v6

@@
 This work advances practical electromagnetic imaging by providing a new, cost-effective paradigm to inverse scattering.
-Equal contribution, random order.
-Correspondence: {ez_cheng, tianchunxun, ou524u}@stu.pku.edu.cn
-{wtzhu, maxiaoxuan, yizhou.wang}@pku.edu.cn
-Project page: https://gomenei.github.io/SingleTX-EISP/
+Code and models are released at https://gomenei.github.io/SingleTX-EISP/.
+Figure 1: Comparison between our method and the previous state-of-the-art.
+Left: Img-Interiors [27] requires case-specific optimization to reconstruct the permittivity.
+In contrast, our method is a data-driven framework that operates in an end-to-end, feed-forward manner for solving inverse scattering.
+Right: Our method yields more accurate reconstructions than Img-Interiors [27].
+It remains robust even with a single transmitter and achieves real-time inference with over $70,000\times$ speed-up.
 1 Introduction
-Electromagnetic waves can penetrate object surfaces, making them essential for non-invasive imaging (Geng et al., 2024; O’Loughlin et al., 2018).
-At the core of electromagnetic imaging lies the Electromagnetic Inverse Scattering Problems (EISP), which seeks to reconstruct an object’s relative permittivity from measured scattered electromagnetic field (Nikolova, 2011).
-By solving EISP, we can accurately recover internal structures without physical intrusion (Song et al., 2005), enabling a range of scientific and industrial applications, such as safer and more cost-effective alternatives to X-rays and MRI scans (Bevacqua et al., 2021; O’Loughlin et al., 2018; Nikolova, 2011).
+Electromagnetic waves can penetrate object surfaces, making them essential for non-invasive imaging [17, 30].
+At the core of electromagnetic imaging lies the Electromagnetic Inverse Scattering Problems (EISP), which seeks to reconstruct an object’s relative permittivity from measured scattered electromagnetic field [29].
+By solving EISP, we can accurately recover internal structures without physical intrusion [38], enabling a range of scientific and industrial applications, such as safer and more cost-effective alternatives to X-rays and MRI scans [5, 30, 29].
 Typically, EISP necessitate a large number of transmitters and receivers to acquire sufficient measurement data.
-This requirement, however, leads to increased operational time and higher costs, thereby limiting the practical applicability of electromagnetic imaging techniques (Leigsnering et al., 2011).
-In contrast, reducing the number of transmitters offers significant advantages, including lower costs and easier deployment in constrained environments.(Baraniuk and Steeghs, 2007; Anitori et al., 2010)
-Figure 1: Comparison between our method and the previous state-of-the-art.
-Left: Img-Interiors (Luo et al., 2024) requires case-specific optimization to reconstruct the permittivity.
-In contrast, our method is a data-driven framework that operates in an end-to-end, feed-forward manner for solving inverse scattering.
-Right: Our method yields more accurate reconstructions than Img-Interiors (Luo et al., 2024).
-It remains robust even with a single transmitter and achieves real-time inference with over $70,000\times$ speed-up.
-However, the inherent ill-posed nature of EISP poses significant challenges to accurate reconstruction (Pan et al., 2011; Chen, 2018; Li et al., 2019; Zhong et al., 2016; Luo et al., 2024), particularly when only a limited number of transmitters are available.
+This requirement, however, leads to increased operational time and higher costs, thereby limiting the practical applicability of electromagnetic imaging techniques [22].
+In contrast, reducing the number of transmitters offers significant advantages, including lower costs and easier deployment in constrained environments.[3, 1, 49]
+However, the inherent ill-posed nature of EISP poses significant challenges to accurate reconstruction [32, 7, 23, 56, 27, 51, 50, 52, 42], particularly when only a limited number of transmitters are available.
 The scarcity of transmitters leads to an insufficient amount of measured data, which fails to capture adequate physical information for stable inversion.
-As a result, approaches relying solely on physical mechanisms(Slaney et al., 1984; Belkebir et al., 2005; Chen, 2009; Zhong and Chen, 2011) often fail to achieve accurate reconstruction.
-Conventional numerical methods such as backpropagation (BP) (Belkebir et al., 2005), generally fail to produce reliable reconstructions under such limited-data conditions.
-Recent machine learning-based approaches like PGAN (Song et al., 2021) and Physics-Net (Liu et al., 2022) often start with an initial solution derived from numerical methods, i.e., BP, and frame the problem as an image-to-image translation task.
+As a result, approaches relying solely on physical mechanisms[37, 4, 6, 55] often fail to achieve accurate reconstruction.
+Conventional numerical methods such as backpropagation (BP) [4], generally fail to produce reliable reconstructions under such limited-data conditions.
+Recent machine learning-based approaches like PGAN [39] and Physics-Net [26] often start with an initial solution derived from numerical methods, i.e., BP, and frame the problem as an image-to-image translation task.
 With only a limited number of transmitters available, reliance on BP becomes a critical bottleneck.
 When BP fails, these methods are unable to correct its errors, as they are not fully end-to-end, ultimately leading to inaccurate reconstructions.
-The most recent method Img-Interiors (Luo et al., 2024) integrates physical mechanisms into neural networks and performs case-by-case optimization.
+The most recent method Img-Interiors [27] integrates physical mechanisms into neural networks and performs case-by-case optimization.
 However, in limited-transmitter scenarios, even after optimization has converged, the resulting reconstructions may still diverge substantially from the ground truth (Fig.˜2), underscoring the intrinsic ambiguity of the inverse problem.
 To address these limitations, we propose a fully end-to-end and data-driven framework that predicts the relative permittivity of scatterers from measured fields, leveraging data distribution priors to compensate for the scarcity of observational data.
+Unlike generative methods where data distribution priors typically refer to explicit, decoupled modules (e.g., denoisers modeling $p(x)$) [21, 10], our method learns an end-to-end mapping that implicitly leverages data statistics to resolve the inherent ill-posedness of electromagnetic inverse scattering problems.
 Specifically, our model takes the measured fields and the spatial coordinate of a position as input and directly predicts the relative permittivity at that location using Multilayer Perceptron (MLP)s, and is trained in a fully end-to-end manner against the ground-truth data.
@@
 2 Related Work
-2.1 Electromagnetic Inverse Scattering Problems (EISP)
+2.1 Electromagnetic Inverse Scattering Problems
 Solving EISP is to determine the relative permittivity of the scatterers based on the scattered field measured by the receivers, thereby obtaining internal imaging of the object.
-The primary challenges of EISP arise from its nonlinearity, ill-posedness, and errors introduced by the discretization (Pan et al., 2011; Chen, 2018; Li et al., 2019; Zhong et al., 2016; Luo et al., 2024).
-Traditional methods for solving EISP can be categorized into non-iterative (Slaney et al., 1984; Devaney, 1981; Habashy et al., 1993; Belkebir et al., 2005) and iterative (Chen, 2009; Zhong and Chen, 2011; Xu et al., 2017; Habashy et al., 1994; van den Berg et al., 1999) approaches.
-Non-iterative methods, such as the Born approximation (Slaney et al., 1984), the Eytov approximation (Devaney, 1981; Habashy et al., 1993), and the BP method (Belkebir et al., 2005), solve nonlinear equations through linear approximations, which inevitably lead to poor quality of the results.
-For better reconstruction quality, iterative methods (Zhong and Chen, 2009; Chen, 2009; Zhong and Chen, 2011; Xu et al., 2017; Habashy et al., 1994; van den Berg et al., 1999; Gao et al., 2015) such as 2-fold Subspace Optimization Method (SOM) (Zhong and Chen, 2009) and Gs SOM (Chen, 2009) are proposed.
-To further overcome the ill-posedness of EISP, diverse regularization approaches and prior information have been widely applied (Oliveri et al., 2017; Shen et al., 2014; Liu et al., 2018; Anselmi et al., 2018).
-However, all of these methods are not generalizable and can be time-consuming because of the iterative schemes (Liu et al., 2022).
+The primary challenges of EISP arise from its nonlinearity, ill-posedness, and errors introduced by the discretization [32, 7, 23, 56, 27].
+Traditional methods for solving EISP can be categorized into non-iterative [37, 13, 18, 4] and iterative [6, 55, 48, 19, 41] approaches.
+Non-iterative methods, such as the Born approximation [37], the Rytov approximation [13, 18], and the BP method [4], solve nonlinear equations through linear approximations, which inevitably lead to poor quality of the results.
+For better reconstruction quality, iterative methods [54, 6, 55, 48, 19, 41, 14] such as 2-fold Subspace Optimization Method (SOM) [54] and Gs SOM [6] are proposed.
+To further overcome the ill-posedness of EISP, diverse regularization approaches and prior information have been widely applied [31, 36, 25, 2].
+However, all of these methods are not generalizable and can be time-consuming because of the iterative schemes [26].
+Figure 2: Difficulties that previous methods faced under a single-transmitter setting.
+(a) BP cannot reconstruct the scatterer.
+(b) Physics-Net makes incorrect guesses.
+(c) Although the reconstruction result of Img-Interiors is consistent with the measured field, the reconstructed scatterer itself is completely different from the ground truth.
 2.2 Machine Learning for EISP
-Recent studies shift to leverage neural networks to solve this problem and demonstrate promising results (Geng et al., 2021; Li et al., 2024).
-Some work (Wei and Chen, 2019; Li et al., 2019; Zhang et al., 2020; Xu et al., 2021; Sanghvi et al., 2019; Song et al., 2021; Liu et al., 2022) adopt a two-stage strategy: they use non-iterative methods such as BP (Belkebir et al., 2005) to generate initial estimates, which are then refined using image-to-image neural networks.
-While these approaches offer a degree of generalization, they are not end-to-end and remain dependent on BP initialization (Belkebir et al., 2005), which becomes their bottleneck.
+Recent studies shift to leveraging neural networks to solve this problem and demonstrate promising results [16, 24].
+Some work [45, 23, 53, 47, 35, 39, 26] adopt a two-stage strategy: they use non-iterative methods such as BP [4] to generate initial estimates, which are then refined using image-to-image neural networks.
+While these approaches offer a degree of generalization, they are not end-to-end and remain dependent on BP initialization [4], which becomes their bottleneck.
 When physical data are too insufficient to reconstruct the scatterer, especially under single-transmitter settings, these approaches tend to “hallucinate” outputs according to unreliable initialization rather than predict the scatterer based on measured field (see Fig.˜5).
-A more recent approach, Img-Interiors (Luo et al., 2024), integrates scattering mechanisms into the network architecture and achieves accurate reconstructions.
-However, it requires case-specific optimization, limiting generalization and making it vulnerable to local minima, often leading to failure in complex settings (see Figs.˜4 and 6).
+A more recent approach, Img-Interiors [27], integrates scattering mechanisms into the network architecture and achieves accurate reconstructions.
+However, it requires case-specific optimization, limiting generalization and making it vulnerable to local minima, often leading to failure in complex settings (see Figs.˜4 and 7).
 Moreover, it fails under a single transmitter setting even when the optimization may have already converged because of ambiguity.
@@
 EISP is the inverse problem of the forward process.
-That is, for an unknown scatterer, we apply certain incident field $\mathbf{E}^{\text{i}}$ to it, and measure the scattered field $\mathbf{E}^{\text{s}}$ as our input.
+That is, for an unknown scatterer, we use transmitters to apply certain incident field $\mathbf{E}^{\text{i}}$ to it, and measure the scattered field $\mathbf{E}^{\text{s}}$ as our input via receivers.
 Our goal is to reconstruct the relative permittivity $\bm{\epsilon}_{r}$ throughout the scatterer.
-For a detailed background introduction of the physical model of EISP, please refer to our supplementary material (Appx.˜B).
+For a detailed introduction of EISP, please refer to our supplementary material (Appx.˜B).
 Specifically, the incident field $\mathbf{E}^{\text{i}}$ excites the induced current $\mathbf{J}$.
-Using the method of moments (Peterson et al., 1998), the total field $\mathbf{E}^{\text{t}}$ for a given transmitter can be expressed as (Colton and Kress, 2013):
+Using the method of moments [33], the total field $\mathbf{E}^{\text{t}}$ for a given transmitter can be expressed as [8]:
 $\mathbf{E}^{\text{t}}=\mathbf{E}^{\text{i}}+\mathbf{G}^{\text{d}}\cdot\mathbf{J},$ (1)
@@
 Since $N_{r}\ll M^{2}$ in practice, reconstructing the induced current $\mathbf{J}$ from the scattered field $\mathbf{E}^{\text{s}}$ is ill-posed.
-Figure 2: Difficulties that previous methods faced under a single-transmitter setting.
-(a) BP cannot reconstruct the scatterer.
-(b) Physics-Net makes incorrect guesses.
-(c) Although the reconstuction result of Img-Interiors is consistent with the measured field, the reconstructed scatterer itself is completely different from the ground truth.
+Figure 3: Overview of our method.
+Our pipeline is built around a MLP that serves as the inverse solver.
+Given the scattered field measurements $\mathbf{E}^{\text{s}}$ from all transmitters and receivers, along with a spatial query $\bf{x}$}, the MLP directly predicts the relative permittivity $\hat{\bm{\epsilon}_{r}}({\bf{x}})$.
+To enhance spatial expressiveness, we apply positional encoding $\gamma({\bf{x}})$ to the query position.
+During training, dashed lines indicate the supervision signals applied.
 Reduction of measured data.
@@
 We divide previous work into three categories and systematically analyze the difficulties they faced under this setting.
-(a) Conventional numerical approaches, such as BP (Belkebir et al., 2005), employ linear approximations, which limit their reconstruction quality.
+(a) Conventional numerical approaches, such as BP [4], employ linear approximations, which limit their reconstruction quality.
 As shown in Fig.˜2, BP cannot even reconstruct a rough shape of the scatterer.
-(b) Machine learning methods based on conventional numerical approaches Song et al.
-(2021); Liu et al.
-(2022), such as Physics-Net.
+(b) Machine learning methods based on conventional numerical approaches [39, 26], such as Physics-Net.
 Although Physics-Net can leverage data-driven training to compensate for missing physical information, its strong dependency on BP initialization becomes a critical bottleneck.
 When BP fails, the model cannot correct the error of BP because it is not fully end-to-end, resulting inaccurate reconstructions, as shown in Fig.˜2.
-(c) Machine learning methods based on implicit functions, such as Img-Interiors (Luo et al., 2024).
+(c) Machine learning methods based on implicit functions, such as Img-Interiors [27].
 Img-Interiors reconstructs a scatterer through case-by-case optimization.
@@
 This approach effectively learns the mapping between scattered field $\mathbf{E}^{\text{s}}$ and relative permittivity $\bm{\epsilon}_{r}$ through training on diverse scattering scenarios, thereby incorporating essential data distribution priors to compensate for the lack of physical information caused by insufficient measurements.
-In the following, we detail our model architecture (Section˜4.2), and the training losses (Section˜4.3).
+In the following, we detail our model architecture (Sec.˜4.2), and the training losses (Sec.˜4.3).
 4.2 Model Architecture
@@
 In the multiple transmitter configuration, $\mathbf{E}^{\text{s}}$ is constructed by combining the complex measurement data from all $N$ transmitters, resulting in a real-valued vector of dimension $2N\cdot N_{r}$ that represents the wave propagation and scattering behavior under diverse illumination conditions provided by transmitters at different locations.
-Figure 3: Overview of our method.
-Our pipeline is built around a MLP that serves as the inverse solver.
-Given the scattered field measurements $\mathbf{E}^{\text{s}}$ from all transmitters and receivers, along with a spatial query $\bf{x}$}, the MLP directly predicts thee relative permittivity $\hat{\bm{\epsilon}_{r}}({\bf{x}})$.
-To enhance spatial expressiveness, we apply positional encoding $\gamma({\bf{x}})$ to the query position.
-During training, dashed lines indicate the supervision signals applied.
 To enhance the model’s capacity to represent high-frequency features, we apply positional encoding to the spatial coordinates $\mathbf{x}_{\text{i}}$, mapping them into a higher-dimensional Fourier feature space using the encoding function: $\gamma(x)=[\sin(x),\cos(x),\ldots,\sin(2^{\Omega-1}x),\cos(2^{\Omega-1}x)]$, where the hyperparameter $\Omega$ controls the spectral bandwidth.
-The complete relative permittivity distribution $\hat{\bm{\epsilon}_{r}}$ is reconstructed by sampling the MLP at all grid points $\{\mathbf{x}_{\text{i}}\}_{i=1}^{M^{2}}$: $\hat{\bm{\epsilon}_{r}}=\{F_{\theta}(\mathbf{E}^{\text{s}},\gamma(\mathbf{x}_{\text{i}})\}_{\text{i}=1}^{M^{2}}.$
+The complete relative permittivity distribution $\hat{\bm{\epsilon}_{r}}$ is reconstructed by sampling the MLP at all grid points $\{\mathbf{x}_{\text{i}}\}_{i=1}^{M^{2}}$: $\hat{\bm{\epsilon}_{r}}=\{F_{\theta}(\mathbf{E}^{\text{s}},\gamma(\mathbf{x}_{\text{i}}))\}_{\text{i}=1}^{M^{2}}.$
 4.3 Training
@@
 This simplified loss function ensures stable and efficient training.
-5 Experiments
-5.1 Setup
-Datasets.
-We train and test our method on standard benchmarks used for EISP following previous work (Wei and Chen, 2019; Song et al., 2021; Liu et al., 2022).
-To enhance training efficiency and model robustness, datasets that share identical transmitter and receiver configurations are combined into a unified training set.
-1) Synthetic Circular-cylinder dataset (Circular(Luo et al., 2024) is synthetically generated comprising images of cylinders with random relative radius, number, and location and permittivity.
-2) Synthetic MNIST dataset (MNIST(Deng, 2012) contains grayscale images of handwritten digits.
-For the two synthetic datasets, we follow previous work(Luo et al., 2024; Deng, 2012), we evaluate two levels of noise: 5% and 30%.
-3) Real-world Institut Fresnel’s database (IF(Geffrin et al., 2005) contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel.
-4) Synthetic 3D MNIST dataset (3D MNIST(de la Iglesia Castro, ) contains 3D data of handwritten digits.
-5) 3D ShapeNet dataset (3D ShapeNet(Wu et al., 2015) contains 3D data of various shapes.
-For more details about datasets, please refer to Appx.˜C.
 Table 1: Quantitative comparison results with SOTA methods.
@@
 MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
-  Number of Transmitters: $N=16$ $N=8/18$
-BP (2005) 0.177 0.719 16.43 0.178 0.716 16.38 0.052 0.905 27.41 0.053 0.904 27.42 0.190 0.779 16.19
-2-fold SOM (2009) 0.154 0.757 20.93 0.156 0.738 20.84 0.031 0.917 32.23 0.038 0.889 30.63 - - -
-Gs SOM (2009) 0.072 0.923 28.31 0.081 0.901 27.13 0.023 0.946 35.40 0.024 0.937 34.89 0.184 0.790 17.00
-BPS (2019) 0.093 0.909 25.00 0.105 0.891 23.90 0.027 0.963 33.00 0.029 0.956 32.42 0.310 0.664 17.05
-PGAN (2021) 0.084 0.916 25.80 0.091 0.910 25.31 0.026 0.966 35.56 0.032 0.947 33.91 0.121 0.926 24.78
-Physics-Net (2022) 0.075 0.932 26.17 0.093 0.906 24.58 0.027 0.934 32.72 0.030 0.927 32.08 0.170 0.788 18.48
-Img-Interiors (2024) 0.200 0.863 26.41 0.336 0.760 19.01 0.036 0.947 35.05 0.047 0.932 32.62 0.153 0.837 23.26
+[…]
+BP [4] 0.177 0.719 16.43 0.178 0.716 16.38 0.052 0.905 27.41 0.053 0.904 27.42 0.190 0.779 16.19
+2-fold SOM [54] 0.154 0.757 20.93 0.156 0.738 20.84 0.031 0.917 32.23 0.038 0.889 30.63 - - -
+Gs SOM [6] 0.072 0.923 28.31 0.081 0.901 27.13 0.023 0.946 35.40 0.024 0.937 34.89 0.184 0.790 17.00
+BPS [7, 45] 0.093 0.909 25.00 0.105 0.891 23.90 0.027 0.963 33.00 0.029 0.956 32.42 0.310 0.664 17.05
+PGAN [39] 0.084 0.916 25.80 0.091 0.910 25.31 0.026 0.966 35.56 0.032 0.947 33.91 0.121 0.926 24.78
+Physics-Net [26] 0.075 0.932 26.17 0.093 0.906 24.58 0.027 0.934 32.72 0.030 0.927 32.08 0.170 0.788 18.48
+Two-Step [51] 0.111 0.835 23.17 0.111 0.835 22.91 0.059 0.830 25.92 0.091 0.848 23.94 - - -
+Img-Interiors [27] 0.200 0.863 26.41 0.336 0.760 19.01 0.036 0.947 35.05 0.047 0.932 32.62 0.153 0.837 23.26
 Ours 0.039 0.978 32.11 0.050 0.966 29.91 0.020 0.965 36.92 0.024 0.954 35.19 0.094 0.916 24.89
-  Number of Transmitters: $N=1$
-BP (2005) 0.194 0.698 15.40 0.194 0.696 15.40 0.065 0.892 25.30 0.065 0.892 25.30 0.199 0.770 16.29
-2-fold SOM (2009) 0.432 0.556 12.49 0.828 0.382 9.45 0.060 0.859 26.63 0.157 0.639 20.07 - - -
-Gs SOM (2009) 0.460 0.598 15.31 0.404 0.557 14.91 0.046 0.888 29.62 0.051 0.862 28.77 0.192 0.779 16.66
-BPS (2019) 0.189 0.774 18.75 0.205 0.744 17.97 0.045 0.891 29.29 0.055 0.862 27.68 0.348 0.669 16.18
-PGAN (2021) 0.133 0.867 21.69 0.153 0.830 20.41 0.033 0.932 32.02 0.040 0.914 29.94 0.248 0.680 16.85
-PhysicsNet (2022) 0.137 0.798 19.98 0.152 0.783 19.38 0.055 0.887 26.60 0.056 0.890 26.48 0.175 0.771 17.45
-Img-Interiors (2024) 0.305 0.604 16.06 0.467 0.484 12.47 0.096 0.855 26.19 0.153 0.806 20.90 0.305 0.705 17.34
+[…]
+BP [4] 0.194 0.698 15.40 0.194 0.696 15.40 0.065 0.892 25.30 0.065 0.892 25.30 0.199 0.770 16.29
+2-fold SOM [54] 0.432 0.556 12.49 0.828 0.382 9.45 0.060 0.859 26.63 0.157 0.639 20.07 - - -
+Gs SOM [6] 0.460 0.598 15.31 0.404 0.557 14.91 0.046 0.888 29.62 0.051 0.862 28.77 0.192 0.779 16.66
+BPS [7, 45] 0.189 0.774 18.75 0.205 0.744 17.97 0.045 0.891 29.29 0.055 0.862 27.68 0.348 0.669 16.18
+PGAN [39] 0.133 0.867 21.69 0.153 0.830 20.41 0.033 0.932 32.02 0.040 0.914 29.94 0.248 0.680 16.85
+PhysicsNet [26] 0.137 0.798 19.98 0.152 0.783 19.38 0.055 0.887 26.60 0.056 0.890 26.48 0.175 0.771 17.45
+Two-Step [51] 0.117 0.845 23.34 0.203 0.656 17.62 0.145 0.673 19.45 0.174 0.652 18.41 - - -
+Img-Interiors [27] 0.305 0.604 16.06 0.467 0.484 12.47 0.096 0.855 26.19 0.153 0.806 20.90 0.305 0.705 17.34
 Ours 0.085 0.921 26.09 0.127 0.862 22.56 0.031 0.931 33.18 0.038 0.914 31.38 0.128 0.908 24.19
+5 Experiments
+5.1 Setup
+Datasets.
+We train and test our method on standard benchmarks used for EISP following previous work [45, 39, 26].
+Datasets that share identical transmitter and receiver configurations are combined into a unified training set.
+1) Synthetic Circular [27] is synthetically generated comprising images of cylinders with random relative radius, number, location, and permittivity.
+2) Synthetic MNIST [12] contains grayscale images of handwritten digits.
+For the two synthetic datasets, following previous work[27, 12], we evaluate two levels of noise: 5% and 30%, and the number of receivers $N_{r}=32$.
+3) Real-world Institut Fresnel’s database (IF[15] contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel, where $N_{r}=241$.
+4) Synthetic 3D MNIST dataset (3D MNIST[11] contains 3D data of handwritten digits.
+5) 3D ShapeNet dataset (3D ShapeNet[46] contains 3D data of various shapes.
+$N_{r}=160$ for the two 3D datasets.
+For more details about datasets, please refer to Appx.˜C.
 Baselines and Metrics.
-We maintain the same settings as in previous studies (Wei and Chen, 2019; Song et al., 2021; Sanghvi et al., 2019) to ensure a fair comparison.
-We compare our method with three traditional methods and four deep learning-based approaches: 1) BP (Belkebir et al., 2005): A traditional non-iterative inversion algorithm.
-2) 2-fold SOM (Zhong and Chen, 2009): A traditional iterative minimization scheme by using SVD decomposition.
-3) Gs SOM (Chen, 2009): A traditional subspace-based optimization method by decomposing the operator of Green’s function.
-4) BPS (Wei and Chen, 2019): A CNN-based image translation method with an initial guess from the BP algorithm.
-5) Physics-Net (Liu et al., 2022): A CNN-based approach that incorporates physical phenomena during training.
-6) PGAN (Song et al., 2021): A CNN-based approach using a generative adversarial network.
-7) Img-Interiors (Liu et al., 2022): An implicit approach optimized by forward calculation.
-Following previous work (Liu et al., 2022), we evaluate the quantitative performance of our method using PSNR (Wang and Bovik, 2009), SSIM (Wang et al., 2004), and Relative Root-Mean-Square Error (MSE(Song et al., 2021).
+For a fair comparison, we follow the same setting as in previous work [45, 39, 35] using their official code and train or optimize each method under the same setup as ours  [^1] BPS, Physics-Net, PGAN and Two-Step are trained as they are learning-based methods; all other baselines are optimization-based; we train separate models for each noise level for both our method and other learning-based models.
+.
+Specifically, we compare our method with 3 traditional methods and 4 deep learning-based approaches: 1) BP [4]: A traditional non-iterative inversion algorithm.
+2) 2-fold SOM [54]: A traditional iterative minimization scheme by using SVD decomposition.
+3) Gs SOM [6]: A traditional subspace-based optimization method by decomposing the operator of Green’s function.
+4) BPS [7, 45]: A CNN-based image translation method with an initial guess from the BP algorithm.
+5) Physics-Net [26]: A CNN-based approach that incorporates physical phenomena during training.
+6) Two-Step [51]: A CNN-based approach with two steps.
+7) PGAN [39]: A CNN-based approach using a generative adversarial network.
+8) Img-Interiors [27]: An implicit approach optimized by forward calculation.
+Following previous work [26], we evaluate the quantitative performance of our method using PSNR [44], SSIM [43], and Relative Root-Mean-Square Error (MSE[39].
 5.2 Comparison with SOTAs
@@
 Colors represent the values of the relative permittivity.
+Figure 5: Qualitative comparison under the single-transmitter setting. Results are obtained with $N=1$ transmitter and noise level of 5%. Colors represent the values of the relative permittivity.
 We begin by comparing our method against prior approaches under the multiple-transmitter setting, using both synthetic and real datasets for comprehensive evaluation.
@@
 In addition, we present a qualitative comparison, as shown in Fig.˜4.
-Traditional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) are only capable of recovering the coarse shape of the scatterer.
-BPS (Chen, 2018) produces sharp edges, but the reconstructed shapes are often inaccurate.
-PGAN (Song et al., 2021) achieves accurate shape recovery, yet introduces noticeable background artifacts.
-Img-Interiors (Luo et al., 2024) can generate high-quality reconstructions, but occasionally fails due to local optima, as it is based on an iterative optimization process (see the last row).
+Traditional methods such as BP[4], Gs SOM [6], and 2-fold SOM [6] are only capable of recovering the coarse shape of the scatterer.
+BPS [7, 45] produces sharp edges, but the reconstructed shapes are often inaccurate.
+PGAN [39] achieves accurate shape recovery, yet introduces noticeable background artifacts.
+Img-Interiors [27] can generate high-quality reconstructions, but occasionally fails due to local optima, as it is based on an iterative optimization process (see the last row).
 In contrast, our method produces accurate and clean reconstructions across all cases, demonstrating both visual fidelity and robustness.
 5.2.2 Single Transmitter Evaluation
-Figure 5: Qualitative comparison under the single-transmitter setting. Results are obtained with $N=1$ transmitter and noise level of 5%. Colors represent the values of the relative permittivity.
+Figure 6: Qualitative comparison under single-transmitter setting on real-world IF dataset.
+The results are obtained with $N=1$ transmitter, without noise level.
+Colors represent the values of the relative permittivity.
 Furthermore, we investigate a highly challenging and practically important setting that has been largely underexplored in previous work: performing EISP with a minimal number of transmitters.
@@
 This remarkable performance under such constrained conditions underscores the efficacy of our end-to-end training framework, which successfully encodes and leverages rich data priors to achieve state-of-the-art results across diverse domains.
-To better understand this phenomenon, we present qualitative comparisons in Fig.˜5.
-Traditional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) produce only blurry reconstructions.
-Deep learning-based methods like BPS (Chen, 2018), Physics-Net (Liu et al., 2022), and PGAN (Song et al., 2021) tend to “hallucinate” the digit, resulting in wrong shape on the MNIST dataset.
-Img-Interiors (Luo et al., 2024) fails to capture the fundamental morphology of the scatterer, resulting in structurally inaccurate representations that deviate significantly from the ground truth.
+To better understand this phenomenon, we present qualitative comparisons in Fig.˜5 for synthetic data and Fig.˜6 for real-world IF [15].
+Traditional methods such as BP[4], Gs SOM [6], and 2-fold SOM [6] produce only blurry reconstructions.
+Deep learning-based methods like BPS [7, 45], Physics-Net [26], and PGAN [39] tend to “hallucinate” the digit, resulting in wrong shape on the MNIST dataset.
+Img-Interiors [27] fails to capture the fundamental morphology of the scatterer, resulting in structurally inaccurate representations that deviate significantly from the ground truth.
 Among all the methods, only ours can still produce reasonably accurate reconstructions of the relative permittivity under such an extreme condition.
@@
 The quantitative results presented in Tab.˜3 demonstrate that our model exhibits smooth and gradual performance degradation as the noise level increases, maintaining excellent reconstruction capability even under strong noise interference as high as 30%.
-Qualitative visualizations in Fig.˜6 show that most baseline methods exhibit noticeable artifacts or even complete failure under severe noise conditions, while our method remains robust and preserves the essential structure of the target.
+Qualitative visualizations in Fig.˜7 show that most baseline methods exhibit noticeable artifacts or even complete failure under severe noise conditions, while our method remains robust and preserves the essential structure of the target.
 Table 2: Ablation study of noise levels effects on MNIST under the multiple-transmitter setting.
@@
 Table 3: Ablation study on training data size under the mutiple-transmitter setting. Noise levels ($5\%$ and $30\%$) in parentheses.
- Data Size MNIST (5%) MNIST (30%)
+ Data MNIST (5%) MNIST (30%)
 […]
@@
 Our method can be naturally extended to 3D scenarios.
-Following previous work (Luo et al., 2024), we employ the Synthetic 3D MNIST (de la Iglesia Castro, ) and extend to 3D ShapeNet (Wu et al., 2015) for training and testing.
-For evaluation, we adopt 3D versions of the MSE (Song et al., 2021) and Intersection over Union (IoU) as our metrics.
+We use the same MLP architecture, with the input dimension extended to 3.
+Following previous work [27], we employ the Synthetic 3D MNIST [11] and extend to 3D ShapeNet [46] for training and testing.
+For evaluation, we adopt 3D versions of the MSE [39] and Intersection over Union (IoU) as our metrics.
 Further details on the datasets are provided in Appx.˜C.
@@
 Fig.˜1 provides a comprehensive comparison of reconstruction quality and runtime between the two methods for both $N=1$ and $N=6$ configurations.
-Fig.˜7 and Fig.˜8 provide visual comparisons of additional 3D reconstruction results on the 3DMNIST and 3DShapeNet datasets.
+Fig.˜8 and Fig.˜9 provide visual comparisons of additional 3D reconstruction results on the 3D MNIST and 3D ShapeNet datasets.
 These results show that our method maintains robustness and generalizes effectively across geometrically complex 3D structures, representing significant progress towards practical applications.
@@
 Limitations.
-While our method effectively handles sparse transmitter settings, it cannot deal with different locations of receivers or transmitters, which remains an important direction for future work.
+While our method effectively handles sparse transmitter settings, it struggles to reconstruct thin structures (see the rightmost block in Fig.˜9 for a typical failure case) and cannot accommodate varying receiver or transmitter locations.
+These limitations challenge deployment in real-world scenarios, where sensor layouts and environmental complexity vary.
+Addressing fine-structure reconstruction and flexible sensor configurations remains an important direction for future work.
+Acknowledgments
+This work was supported by NSFC-6247070125, Ningbo Grants 2025Z038 and 2025Z059, and the High Performance Computing Center at Eastern Institute of Technology and Ningbo Institute of Digital Twin.
 Appendix A The Use of Large Language Models (LLMs)
@@
 So, it interacts with the sum of the incident field and the scattered field, aka the total field $E^{\text{t}}$.
-The total field can be described by Lippmann-Schwinger equation (Colton and Kress, 2013) as follows:
+The total field can be described by Lippmann-Schwinger equation [8] as follows:
 $E^{\text{t}}({\bf{x}})={E}^{\text{i}}({\bf{x}})+k_{0}^{2}\int_{\mathcal{D}}g({\bf{x}},{\bf{x^{\prime}}})J({\bf{x^{\prime}}})d{\bf{x^{\prime}}},{\bf{x}}\in\mathcal{D},$ (4)
@@
 where ${\bf{x}}\in S$ indicates that the scattered field is measured by the receivers at surface $S$ around $\mathcal{D}$.
-Since digital analysis only applies to discrete variables (Mitra, 2001; Taflove et al., 2005), we discretize equations Eq.˜4, Eq.˜5, and Eq.˜6.
-The region of interest $\mathcal{D}$ is discretized into $M\times M$ square subunits, and we use the method of moment (Peterson et al., 1998) to obtain the discrete scattered field $\mathbf{E}^{\text{s}}$ [^1] We use the bold letters to represent discrete variables..
+Since digital analysis only applies to discrete variables [28, 40], we discretize equations Eq.˜4, Eq.˜5, and Eq.˜6.
+The region of interest $\mathcal{D}$ is discretized into $M\times M$ square subunits, and we use the method of moment [33] to obtain the discrete scattered field $\mathbf{E}^{\text{s}}$ [^2] We use the bold letters to represent discrete variables..
 The discrete version of Eq.˜4 is as follows:
@@
 To enhance model robustness and training efficiency, datasets that share identical measurement settings are pooled to form consolidated training sets.
-We strictly follow the established rules in EISP (Wei and Chen, 2019; Song et al., 2021; Liu et al., 2022) to set the dataset.
+We strictly follow the established rules in EISP [45, 39, 26] to set the dataset.
 We train and test our model on standard benchmarks used for EISP.
-1) Circular (Luo et al., 2024) is synthetically generated comprising images of cylinders with random relative radius, number, location, and permittivity between $1$ and $1.5$.
+1) Circular [27] is synthetically generated comprising images of cylinders with random relative radius, number, location, and permittivity between $1$ and $1.5$.
 $10k$ images are generated for training purposes, and $1.2k$ images for testing.
-2) MNIST (Deng, 2012) contains grayscale images of handwritten digits.
-Similar to previous settings (Wei and Chen, 2019; Zhou et al., 2022), we use them to synthesize scatterers with relative permittivity values between $2$ and $2.5$ according to their corresponding pixel values.
+2) MNIST [12] contains grayscale images of handwritten digits.
+Similar to previous settings [45, 57], we use them to synthesize scatterers with relative permittivity values between $2$ and $2.5$ according to their corresponding pixel values.
 The entire MNIST training set containing $60k$ images is used for training purposes, while $1.2k$ images from the MNIST test set are randomly selected for testing.
@@
 All transmitters and receivers are equally placed.
+For single-transmitter settings, the transmitter is positioned at the maximum x-coordinate at (3, 0).
 (b) For IF dataset, we set up $N=8$ or $18$ transmitters and $N_{r}=241$ receivers.
@@
 For all datasets, the green square represents the region of interest $\mathcal{D}$.
-Following previous work (Chen, 2018; Zhou et al., 2022) to generate the above two synthetic datasets, we set operating frequency $f=400$ MHz.
+Following previous work [7, 57] to generate the above two synthetic datasets, we set operating frequency $f=400$ MHz.
 The region of interest is a square with the size of $2m\times 2m$.
@@
 The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in LABEL:fig:schematic_synthetic.
-The data are generated numerically using the method of moments (Peterson et al., 1998) with a $224\times 224$ grid mesh to avoid inverse crime (Colton et al., 1998).
+The data are generated numerically using the method of moments [33] with a $224\times 224$ grid mesh to avoid inverse crime [9].
 To simulate the noise in actual measurement, we add a $5\%$ level of noise to the scattered field $\mathbf{E}^{\text{s}}$ for regular setting.
-3) Real-world IF dataset (Geffrin et al., 2005) contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel.
+3) Real-world IF dataset [15] contains three different dielectric scenarios, namely FoamDielExt, FoamDielInt, and FoamTwinDiel.
 $N=8$ for FoamDielExt and FoamDielInt, $N=18$ for FoamTwinDiel, and $N_{r}=241$ for all the cases.
@@
 Figure 11: Positions of transmitters and receivers for 3D MNIST dataset. There are $N=40$ transmitters and $N_{r}=160$ receivers. The green cube represents the region of interest $\mathcal{D}$.
-4) Synthetic 3D MNIST (de la Iglesia Castro, ) contains 3D images of handwritten digits.
-5) Synthetic 3D ShapeNet (Wu et al., 2015) contains 3D images of various shapes.
+4) Synthetic 3D MNIST [11] contains 3D images of handwritten digits.
+5) Synthetic 3D ShapeNet [46] contains 3D images of various shapes.
 We use these two datasets to synthesize scatterers with relative permittivity value of $2$.
@@
 For the positions of receivers, the azimuthal angle ranges from $0^{\circ}$ to $348.75^{\circ}$ with the step of $11.25^{\circ}$, and the polar angle ranges from $30^{\circ}$ to $150^{\circ}$ with the step of $30^{\circ}$.
-The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in Fig.˜10.
+The schematic diagram of the locations of the transmitters and receivers around the region of interest is shown in Fig.˜11.
 The entire 3D MNIST dataset is used, including $5k$ images for training purposes and $1k$ images for testing.
@@
 C.2 Metrics
-Following previous work (Liu et al., 2022), we evaluate the quantitative performance of our method using PSNR (Wang and Bovik, 2009), SSIM (Wang et al., 2004), Relative Root-Mean-Square Error (MSE) (Song et al., 2021) and IoU.
+Following previous work [26], we evaluate the quantitative performance of our method using PSNR [44], SSIM [43], Relative Root-Mean-Square Error (MSE) [39] and IoU.
 For PSNR, SSIM and IoU, a higher value indicates better performance.
@@
 Appendix D Additional Experimental Results
-This section provides additional experimental results, including statistical validation (Section˜D.1) and extended qualitative comparisons (Section˜D.2).
-The latter includes more visual comparisons with baseline methods, as well as supplementary ablation studies on noise robustness and training data size.
+This section provides additional experimental results, including statistical validation (Sec.˜D.1), extended qualitative comparisons (Sec.˜D.2) and additional ablation studies on loss functions (Sec.˜D.3).
+The qualitative comparisons include more visual comparisons with baseline methods, while the ablations further analyze noise robustness, training data size, and the impact of different training losses.
 D.1 Experimental Statistical Significance
@@
 MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
-  Number of Transmitters: $N=16$
+[…]
 mean 0.039 0.978 32.11 0.050 0.966 29.91 0.020 0.965 36.92 0.024 0.954 35.19
 std $2.8\times 10^{-5}$ $5.2\times 10^{-5}$ $6.2\times 10^{-3}$ $2.1\times 10^{-4}$ $2.4\times 10^{-4}$ $1.7\times 10^{-2}$ $1.6\times 10^{-5}$ $4.8\times 10^{-5}$ $7.6\times 10^{-3}$ $7.1\times 10^{-5}$ $3.3\times 10^{-4}$ $1.4\times 10^{-2}$
-  Number of Transmitters: $N=1$
+[…]
 mean 0.085 0.921 26.09 0.127 0.862 22.56 0.031 0.931 33.18 0.038 0.914 31.38
@@
 Qualitative Comparison.
-We present more qualitative comparison on Circular dataset (Luo et al., 2024) and MNIST dataset (Deng, 2012) under settings with different transmitter numbers $N$ and noise levels, as shown in Fig.˜11 to Fig.˜16.
-We also present qualitative comparison on IF (Geffrin et al., 2005) dataset under the single-transmitter setting, as shown in Fig.˜17.
-Our method achieves comparable or superior performance to SOTA methods in most cases under multiple-transmitter settings, as Fig.˜11, Fig.˜12, Fig.˜15, and Fig.˜16 indicate.
-As shown in Fig.˜12 and Fig.˜16, PGAN (Song et al., 2021) introduces noticeable background artifacts due to the lack of consideration of physics.
-And Fig.˜11 shows that Img-Interiors occasionally fails to converge due to local optima.
-For single-transmitter setting ($N=1$), our method significantly outperforms all previous approaches across all datasets, as shown in Fig.˜13 and Fig.˜14.
-Conventional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) produce only blurry reconstructions.
-Deep learning-based methods such as BPS (Chen, 2018), Physics-Net (Liu et al., 2022), and PGAN (Song et al., 2021) tend to “hallucinate” the digit, leading to wrong reconstruction on the MNIST dataset.
-Img-Interiors (Luo et al., 2024) produces results with structural errors that deviate significantly from the true morphology.
+We present more qualitative comparison on Circular dataset [27] and MNIST dataset [12] under settings with different transmitter numbers $N$ and noise levels, as shown in Fig.˜12 to Fig.˜17.
+Our method achieves comparable or superior performance to SOTA methods in most cases under multiple-transmitter settings, as Fig.˜12, Fig.˜13, Fig.˜16, and Fig.˜17 indicate.
+As shown in Fig.˜13 and Fig.˜17, PGAN [39] introduces noticeable background artifacts due to the lack of consideration of physics.
+And Fig.˜12 shows that Img-Interiors occasionally fails to converge due to local optima.
+For single-transmitter setting ($N=1$), our method significantly outperforms all previous approaches across all datasets, as shown in Fig.˜14 and Fig.˜15.
+Conventional methods such as BP[4], Gs SOM [6], and 2-fold SOM [6] produce only blurry reconstructions.
+Deep learning-based methods such as BPS [7], Physics-Net [26], and PGAN [39] tend to “hallucinate” the digit, leading to wrong reconstruction on the MNIST dataset.
+Img-Interiors [27] produces results with structural errors that deviate significantly from the true morphology.
 In contrast, our method can still produce reasonably accurate reconstructions of the relative permittivity under such a challenging condition.
@@
 3D Reconstruction.
-We present more qualitative comparison on 3D MNIST (de la Iglesia Castro, ) dataset under the single-transmitter setting, as shown in Fig.˜18.
-Our method successfully approximates permittivity reconstruction even in this difficult setting, whereas Img-Interiors (Luo et al., 2024) fails.
+We present more qualitative comparison on 3D MNIST [11] dataset under the single-transmitter setting, as shown in Fig.˜18.
+Our method successfully approximates permittivity reconstruction even in this difficult setting, whereas Img-Interiors [27] fails.
 Similar to the 2D scenario, Img-Interiors suffers from degeneration and can only produce fixed patterns, failing to capture the shape of scatterers.
-Additionally, we evaluate our method on the more complex 3D ShapeNet (Wu et al., 2015) dataset to demonstrate its generalization capability.
+Additionally, we evaluate our method on the more complex 3D ShapeNet [46] dataset to demonstrate its generalization capability.
 As shown in Fig.˜19, our approach can successfully reconstruct various complex structures including airplanes, cars and tables under the same single-transmitter configuration, maintaining accurate permittivity distribution recovery.
@@
 Noise Robustness.
-We present qualitative results across multiple noise levels (ranging from 5% to 30%) on the MNIST (Deng, 2012) dataset, providing visual support for the noise robustness analysis.
+We present qualitative results across multiple noise levels (ranging from 5% to 30%) on the MNIST [12] dataset, providing visual support for the noise robustness analysis.
 As shown in Fig.˜20, our method maintains high visual fidelity at low noise levels.
@@
 Ablation on Training Data Size.
-We present qualitative results across different training data scales (ranging from 25% to 100%) on the MNIST (Deng, 2012) dataset, offering visual insights into the impact of data volume on reconstruction performance.
+We present qualitative results across different training data scales (ranging from 25% to 100%) on the MNIST [12] dataset, offering visual insights into the impact of data volume on reconstruction performance.
 As shown in Fig.˜21 and Fig.˜22 , our model maintains satisfactory reconstruction integrity even with limited training data.
+D.3 Additional Ablation Results
+We experiment with TV loss [34] for smoothness and Perceptual loss [20] for structure in addition to MSE on single-transmitter setting.
+Tab.˜5 reports the quantitative results on the MNIST dataset under different noise levels.
+Overall, the auxiliary losses provide mixed improvements, while MSE alone already provides strong and stable performance, indicating that our model is relatively robust to the choice of loss function.
+Table 5: Ablation of loss functions.
+We compare MSE with auxiliary TV loss [34] and perceptual loss [20] under the single-transmitter setting.
+Results are reported on the MNIST dataset under two noise levels: 5% and 30%.
+Loss MNIST (5%) MNIST (30%)
+MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
+Ours (MSE) 0.085 0.921 26.09 0.127 0.862 22.56
+Ours (w/ TV) 0.084 0.917 26.00 0.124 0.859 22.74
+Ours (w/ Percpt.) 0.084 0.923 26.20 0.130 0.861 22.49
 Figure 12: Qualitative comparison under the multiple-transmitter setting on MNIST dataset.
@@
 Colors represent the values of the relative permittivity.
-Figure 17: Qualitative comparison under single-transmitter setting on IF dataset.
-The results are obtained with $N=1$ transmitter, without noise level.
-Colors represent the values of the relative permittivity.
 Figure 18: Qualitative comparison under the single-transmitter setting for 3D reconstruction on 3D MNIST dataset.